OpenStudy (anonymous):

d/dx m(x) where m(x) is the tangent line to curve x^2+xy-y^2=1 at point (2,3)

OpenStudy (anonymous):

my work: \( 2x+y+xy'-2yy'=0\\ 2x+y=2yy'-xy'\\ 2x+y=y'(2y-x)\\ \dfrac{dy}{dx}=\dfrac{2x+y}{2y-x} \)

OpenStudy (anonymous):

do i just plug the point in now? or do I use y-y1=m(x-x1) somehow

OpenStudy (mathmate):

Since m'(x)=the slope m at (2,3), y'(2) should equal m'(x).

OpenStudy (mathmate):

* y'(2,3)

OpenStudy (anonymous):

I'm not sure I understand what the question is asking? is 7 the answer?

OpenStudy (loser66):

\(\dfrac{d}{dx}(m(x))\) means : after you got m(x), take derivative of it again.

OpenStudy (anonymous):

ohhh

OpenStudy (loser66):

and then plug (2,3) in

OpenStudy (anonymous):

it's been like 2 years since I took a calc class, was helping someone with this question. thank you both

OpenStudy (anonymous):

thank you https://www.desmos.com/calculator/ykoruebu4u

OpenStudy (mathmate):

 The question is asking for the slope of the tangent line m(x), which should be the same as the slope of the curve at (2,3), by definition of the tangent. that is why m'(x)=slope of tangent line = y'(2,3)